OpenStudy (anonymous):

7+9 lnx = 4 Solve and explain please!

OpenStudy (harsimran_hs4):

9 lnx = (4 - 7)= -3 lnx = -1/3 x = e^(-1/3)

OpenStudy (anonymous):

Oy, yup that's it. (: Thankss!